A pentagon has two right angles, two obtuse angles that are congruent, and a fifth angle that measures 40 degrees less than twice the measure of one of the obtuse angles. What is the measure of the fifth angle?
Answer by richard1234(7193)

Suppose the angles are 90, 90, x, x, and 2x-40, where x>90. The sum of the angles is 540 degrees, so

90 + 90 + x + x + (2x-40) = 540

4x + 140 = 540

x = 100

So, the fifth angle is 2(100) - 40 = 160 degrees.